# Vision Transformers on the Edge: A Comprehensive Survey of Model Compression and Acceleration Strategies

### v1

@@ 2.3.2.2 QAT Techniques for Vision Transformer
Algorithm Key point Backbone Dataset Results
@@
PTQ  [89] Similarity-aware quantization for linear layers ranking -aware quantization for self-attention layers Mixed-precision quantization to retain performance ViT-B CIFAR-10 Percentile 6 MP 6 MP 96.83 ($+3.35$)
@@
DeiT-B ImageNet-1k [51] Bit-Split 6 MP 6 MP 74.58 ($+0.54$)
PTQ4ViT  [84] Twin uniform method to reduce the quantization error on activation values & analyse Hessian guided metric to determine the scaling factors of each layer ViT-B ImageNet-1k [51] Base-PTQ 8 8 85.82 ($+0.52$)
@@
APQ-ViT  [85] Solve for extremely low-bit representation; BBC to apply quantization in a blockwise manner to perceive the loss in adjacent layers .& Matthew-effect preserving quantization for the softmax to maintain power-law distribution ViT-B ImageNet-1k [51] PTQ4ViT [84] 6 6 82.21 ($+0.56$)
@@
NoisyQuant  [86] A quantizer-agnostic enhancement for the post-training activation quantization of ViT & adding a fixed Uniform noisy bias to the values being quantized for a given quantizer ViT-B ImageNet-1k [51] PTQ4ViT [84] 6 6 81.90 ($+6.24$)
@@
FQ-Vit  [14] Efficient PTQ method for achieving accurate quantization on LayerNorm inputs with one layerwise quantization scale named PTF. Propose LIS for performing 4-bit quantization on attention maps DeiT-T ImageNet-1k [51] Percentile 8 8 71.61 ($+0.14$)
@@
I-ViT  [87] Performing the entire inference with integer arithmetic & bit-shifting. integer approximations for non-linear operations ViT-B ImageNet-1k [51] I-BERT [88] - - 84.76 ($+1.06$)
@@
RepQ-ViT  [15] Apply channel-wise quantization on post-LayerNorm activations & $\log\sqrt{2}$ for post-softmax activations ViT-B ImageNet-1k [51] APQ-ViT [85] 4 4 83.62 ($+1.41$)
@@
MPTQ-ViT  [92] Introduce SmoothQuant [94] with bias term to solve asymmetric issue & minimize clamping loss ViT-B ImageNet-1k [51] TSPTQ-ViT [98] 6 6 82.70 ($+0.41$)
@@
LRP-QViT  [93] Assigning precision bit for individual layers based on layer’s importance ViT-B ImageNet-1k [51] RepQ-ViT [15] 6 MP 6 MP 83.87 ($+0.25$)
@@
ERQ  [95] Introduced weight quantization error reduction metrics to minimize the weight quantization error ViT-B ImageNet-1k [51] AdaRound [99] 5 5 82.81 ($+0.81$)
@@
Adalog  [96] Proposed adaptive log based non-uniform quantization for post-Softmax & post GELy activations ViT-B ImageNet-1k [51] RepQ-ViT [15] 6 6 84.80($+1.18$)
@@
CLAMP-ViT  [97] Leverage contrastive learning layer-wise evolutionary search for fixed and mixed-precision quantization DeiT-S ImageNet-1k [51] LRP-QViT [93] 6 6 79.43 ($+0.40$)
@@
Leveraging KD in QAT Q-Vit [100] first proposed an information rectification module based on information theory to resolve the convergence issue during joint training of quantization. The authors then proposed distributed guided distillation by taking appropriate activities and utilizing the knowledge from similar matrices in distillation to perform the optimization perfectly. However, Q-ViT lacks other CV tasks, such as object detection. Another recent work, Q-DETR [79], is introduced to solve the information distortion problem. The authors explored the low-bits quantization of DETR and proposed a bi-level optimization framework based on the information bottleneck principle. However, Q-DETR failed to keep the attention activations less than 4 bits and resulted in mixed-precision quantization, which is hardware-inefficient in the current scenario. Both Q-ViT and Q-DETR explored the lightweight version of DETR apart from modifying the MHA. AQ-DETR [101] focused on solving the problem that exists for low bits of DETR in previous studies. The authors introduced an auxiliary query module and layer-by-layer distillation module to reduce the quantization error between quantized attention and full-precision counterpart. All the previously discussed works are heavily dependent on the data. Li et al.  [102] proposed PSAQ-ViT, aiming to achieve a data-free quantization framework by utilizing the property of KD. The authors introduced an adaptive teacher-student strategy enabling cyclic interaction between generated samples and the quantized model under the supervision of the full-precision model, significantly improving accuracy. The framework leverages task- and model-independent prior information, making it universal across various vision tasks such as classification and object detection.
@@
Standalone QAT techniques Although most works utilized the KD in QAT works, there are limited standalone studies without KD. PackQViT [106] proposed activation-aware sub-8-bit QAT techniques for mobile devices. The authors leveraged $\log 2$ quantization or clamping to address the long-tailed distribution and outlier-aware training to handle the channel-wise outliers. Furthermore, The authors utilized int-2n-softmax, int-LayerNorm, and int-GELU to enable integer-only computation. They designed a SIMD-based 4-bit packed multiplier to achieve end-to-end ViT acceleration on mobile devices. Another recent study named QD-BEV [67] explored QAT on BEVFormer [107] by leveraging image and BEV features. The authors identified that applying quantization directly in BEV tasks makes the training unstable, which leads to performance degradation. The authors proposed view-guided distillation to stabilize the QAT by conducting a systematic analysis of quantizing BEV networks. This work will open a new direction for autonomous vehicle research, applying QAT to reduce computational costs.
Table 9: Results of different quantization-aware training techniques proposed for ViTs. C and OD refer to classification tasks and object detection tasks in the CV tasks column, respectively. CMR refers to the Cascade Mask R-CNN model. The results of QD-BEV [67] is the mean average precision (mAP) rather than APbox.
Algorithm Key point CV tasks Dataset Results
@@
Q-ViT  [100] Propose a switchable scale to resolve convergence issue during joint training of quantization scales & bit-widths; limits of ViT quantization to 3-bit C ImageNet-1k [51] LSQ [108] DeiT-B 4 4 83.0 ($+2.1$) -
@@
Q-DETR  [79] Utilized knowledge distillation to improve the representation capacity of the quantized model OD PASCAL VOC [109] LSQ [108] DETR-R50 2 2 - 50.7 ($+8.1$)
SMCA-DETR-R50 2 2 - 50.2 ($+7.9$)
COCO 2017 [103] DETR-R50 4 4 - 39.4 ($+6.1$)
SMCA-DETR-R50 4 4 - 38.3 ($+4.4$)
AQ-DETR  [101] Proposed a QAT technique based on auxiliary queries for DETR OD PASCAL VOC [109] Q-DETR [79] DETR-R50 4 4 - 53.7 ($+3.3$)
Deformable DETR-R50 4 4 - 63.1 ($+2.0$)
COCO 2017  [103] DETR-R50 4 4 - 40.2 ($+2.8$)
Deformable DETR-R50 4 4 - 44.1 ($+3.4$)
PSAQ-ViT V2  [102] Proposed a data-free quantization framework with adaptive teacher-student strategy C & OD ImageNet-1k [51] PSAQ-ViT [110] DeiT-B 8 8 81.5 ($+2.4$) -
@@
COCO 2017 [103] Standard V2 DeiT-S + CMR 8 8 - 44.8 (+.$+.3$)
Swin-S + CMR 8 8 - 50.9 ($+0.6$)
PackQViT  [106] Proposed an 8-bit QAT framework for mobile devices C & OD ImageNet-1k [51] Q-ViT [100] DeiT-B 8 8 82.9 ($+0.5$) -
@@
COCO 2017 [103] PTQ [89] DETR-R50 8 8 - 60.0 ($-3.1$)
QD-BEV  [67] Introduced a view-guided objective to stabilize the QAT training for image features & BEV features OD NuScenes [111] BEVFormer-B-DFQ [112] BEVFormer 8 8 - 40.6 ($+2.2$)
BEVFormer-S-HAWQv3 [113] 8 8 - 40.6 ($+3.0$)
BEVFormer-B-PACT [114] 8 8 - 40.6 ($+3.2$)
2.3.3 Discussion
Tables 7 and 8 provide an overview of PTQ techniques for image classification and object detection on ViT models. A notable trend is the dominance of post-training quantization (PTQ) methods, with fewer studies exploring training-phase QAT. Table 7 highlights the top-1 accuracy improvements achieved by PTQ methods on ViT architectures for the classification tasks, while Table 8 summaries the mean average precision (mAP) and APbox improvement achieved from the baseline techniques for the object detection task. Moreover, Table 9 summarizes QAT techniques for image classification and object detection. Interestingly, most QAT methods for ViTs, such as those leveraging knowledge distillation [100, 79, 101, 102], focus on optimizing quantized models but lack experimental validation on edge devices. In contrast, approaches like PackQViT [106] have experimented with their QAT frameworks on mobile devices, pushing the boundaries of practical deployment. However, these quantization techniques broadly experimented on datasets like ImageNet-1k and COCO 2017, raising questions about their generalizability to specialized domains such as medical imaging or autonomous driving. This gap underscores the need for future research to explore versatile quantization strategies that cater to diverse application areas and resource-constrained edge devices.
@@ 3.1 Software Tools
Deploying deep learning models on heterogeneous platforms demands specialized software tools that bridge the gap between cutting-edge artificial intelligence (AI) research and real-world applications. These tools empower developers to optimize, accelerate, and seamlessly integrate AI models across different hardware architectures. Table 10 illustrates the most popularly used software tools/libraries/engines to deploy the deep learning models on different hardware architectures. The software libraries are divided into three hardware architectures: FPGA, GPU, and CPU. As FPGAs offer highly parallel and reconfigurable hardware capabilities, deploying AI models on FPGAs requires specialized software tools for efficient hardware mapping, optimization, and deployment. Both Vivado Design Suite  [^3] Vivado design suite. Retrieved January 18, 2025, from https://www.amd.com/en/products/software/adaptive-socs-and-fpgas/vivado.html (from Xilinx) and the Quartus Prime Design Software  [^4] Intel Quartus Prime. Retrieved January 18, 2025, from https://www.intel.com/content/www/us/en/products/details/fpga/development-tools/quartus-prime.html (from Intel) offer advanced synthesis, converting high-level languages to hardware description language (HDL) and preoptimized AI accelerators IP cores (such as Xilinx DPU or Intel AI Suite) that help to accelerate the inference task. Vitis AI  [^5] Vitis AI. Retrieved January 18, 2025, from https://www.xilinx.com/products/design-tools/vitis/vitis-ai.html is the software platform for Xilinx FPGA while OpenVINO  [^6] Openvinotoolkit. Retrieved January 18, 2025, from https://github.com/openvinotoolkit/openvino designed for Intel FPGA and other hardware architectures, including GPU and CPU. However, It is possible/likely that each of these environments is highly modified only for their hardware family, which means developing applications on one would make it very difficult to port to the other. Besides those two software tools, Hls4ml  [^7] Hls4ml. Retrieved January 18, 2025, from https://fastmachinelearning.org/hls4ml/index.html and FINN  [^8] FINN. Retrieved January 18, 2025, from https://xilinx.github.io/finn/ designed to explore deep neural network inference on FPGAs efficiently and swiftly. However, both these libraries are still in the experimental phase. In the software tools for GPU, NVIDIA has a wide range of libraries/engines for edge devices. NVIDIA Triton Inference Server  [^9] NVIDIA Triton Inference Server. Retrieved January 18, 2025, from https://github.com/triton-inference-server/server is the most prominent open-source and scalable inference-serving software engine that simplifies the deployment of deep learning models at scale across all NVIDIA GPUs, x86, and Arm CPUs from major frameworks, including TensorFlow, PyTorch, and NVIDIA TensorRT. Additionally, TensorRT  [^10] TensorRT. Retrieved January 18, 2025, from https://developer.nvidia.com/tensorrt uses as a popular inference optimizer and runtime library for NVIDIA GPU-based edge devices. oneDNN 111111oneDNN. Retrieved January 18, 2025, from https://github.com/oneapi-src/oneDNN is a widely used open-source, cross-platform performance tool for deep learning models. It is optimized for Intel processors, graphics, and ARM-based processors and is in the experimental stage for NVIDIA GPU, AMD GPU, and RISC-V processors. Each software tool is mostly designed to optimize performance and efficiency for specific hardware architectures. However, ONNX Runtime  [^12] ONNX Runtime. Retrieved January 18, 2025, from https://onnxruntime.ai/ is one of the few inference engines that supports a wide range of hardware, including CPUs, GPUs, and FPGAs.
@@
Tool Name Type Key features Supported edge devices
FPGA GPU CPU
Xilinx Vivado design suite 3 Toolkit Used for synthesis, simulation, and configuring Xilinx FPGAs
@@
Intel Quartus Prime 4 Toolkit Used for synthesis, place-and-route, and programming.
@@
OpenVINO 6 Toolkit Optimized inference, supports multiple frameworks, Intel hardware-focused
@@
Vitis AI 5 Engine Model compression, xmodel generation, diverse pretrained model
@@
Hls4ml 7 Library High-level synthesis for FPGA, supports various frameworks Yes No No
FINN footreffinn Engine Quantized models, based on Vitis AI, streaming dataflow for inference
@@
TensorRT 10 Engine High-performance AI inference on NVIDIA GPUs No
@@
NVIDIA Triton Inference Server 9 Inference Server Scalable inference serving, multi-model deployment, Provide model analyzer No
@@
oneDNN 11 Library Optimized deep learning performance, cross-platform No Experimental Intel CPU, ARM
ONNX Runtime 12 Engine Open-source inference, hardware-specific execution providers Yes Yes Yes

### v2

@@ 2.3.2.2 QAT Techniques for Vision Transformer
Algorithm Code Key point Backbone Dataset Results
@@
PTQ  [70] Introduce a mixed-precision weights strategy ViT-B CIFAR-10 Percentile 6 MP 6 MP 96.83 ($+3.35$)
@@
DeiT-B
ImageNet-1k
 [45]
Bit-Split 6 MP 6 MP 74.58 ($+0.54$)
PTQ4ViT  [66] Twin-uniform quantization with Hessian-guided calibration ViT-B ImageNet-1k  [45] Base-PTQ 8 8 85.82 ($+0.52$)
@@
APQ-ViT  [88] Propose a Matthew-effect preserving scheme for low bit quantization ViT-B ImageNet-1k  [45] PTQ4ViT  [66] 6 6 82.21 ($+0.56$)
@@
NoisyQuant  [92] Add a fixed noisy factor to the date distribution ViT-B ImageNet-1k  [45] PTQ4ViT  [66] 6 6 81.90 ($+6.24$)
@@
FQ-Vit  [14] Power-of-two for LayerNorm & replace Softmax with Log-Int-Softmax DeiT-T ImageNet-1k  [45] Percentile 8 8 71.61 ($+0.14$)
@@
I-ViT  [93] Integer-only inference ViT-B ImageNet-1k  [45] I-BERT  [94] - - 84.76 ($+1.06$)
@@
RepQ-ViT  [15] $\log_{\sqrt{2}}$ reparametrized to $\log_{2}$ during quantization & inference ViT-B ImageNet-1k  [45] APQ-ViT  [88] 4 4 83.62 ($+1.41$)
@@
MPTQ-ViT  [97] Introduce SmoothQuant [98] with bias term ViT-B ImageNet-1k  [45] TSPTQ-ViT  [102] 6 6 82.70 ($+0.41$)
@@
LRP-QViT  [80] Importance-based layer bit precision ViT-B ImageNet-1k  [45] RepQ-ViT  [15] 6 MP 6 MP 83.87 ($+0.25$)
@@
ERQ  [99] Introduced weight quantization error reduction metrics ViT-B ImageNet-1k  [45] AdaRound  [103] 5 5 82.81 ($+0.81$)
@@
Adalog  [100] Proposed adaptive log-based quantization for post-Softmax & post GELU activations ViT-B ImageNet-1k  [45] RepQ-ViT  [15] 6 6 84.80($+1.18$)
@@
CLAMP-ViT  [101] Evolutionary search for optimal quantization parameters DeiT-S ImageNet-1k  [45] LRP-QViT  [80] 6 6 79.43 ($+0.40$)
@@
Leveraging KD in QAT Q-Vit [73] first proposed an information rectification module based on information theory to resolve the convergence issue during joint training of quantization. The authors then proposed distributed guided distillation by taking appropriate activities and utilizing the knowledge from similar matrices in distillation to perform the optimization perfectly. However, Q-ViT lacks other CV tasks, such as object detection. Another recent work, Q-DETR [84], is introduced to solve the information distortion problem. The authors explored the low-bit quantization of DETR and proposed a bi-level optimization framework based on the information bottleneck principle. However, Q-DETR failed to keep the attention activations less than 4 bits and resulted in mixed-precision quantization, which is hardware-inefficient in the current scenario. Both Q-ViT and Q-DETR explored the lightweight version of DETR apart from modifying the MSA. AQ-DETR [104] focused on solving the problem that exists for low bits of DETR in previous studies. The authors introduced an auxiliary query module and a layer-by-layer distillation module to reduce the quantization error between the quantized attention and its full-precision counterpart. All the previously discussed works are heavily dependent on the data. Li et al.  [105] proposed PSAQ-ViT, aiming to achieve a data-free quantization framework by utilizing the property of KD. The authors introduced an adaptive teacher-student strategy enabling cyclic interaction between generated samples and the quantized model under the supervision of the full-precision model, significantly improving accuracy. The framework leverages task- and model-independent prior information, making it universal across various vision tasks such as classification and object detection.
@@
Standalone QAT techniques Although most works utilized the KD in QAT works, there are limited standalone studies without KD. PackQViT [109] proposed activation-aware sub-8-bit QAT techniques for mobile devices. The authors leveraged $\log_{2}$ quantization or clamping to address the long-tailed distribution and outlier-aware training to handle the channel-wise outliers. Furthermore, the authors utilized int-2n-softmax, int-LayerNorm, and int-GELU to enable integer-only computation. They designed a SIMD-based 4-bit packed multiplier to achieve end-to-end ViT acceleration on mobile devices. Another recent study named QD-BEV [72] explored QAT on BEVFormer [110] by leveraging image and BEV features. The authors identified that applying quantization directly in BEV tasks makes the training unstable, which leads to performance degradation. The authors proposed view-guided distillation to stabilize the QAT by conducting a systematic analysis of quantizing BEV networks. This work will open a new direction for autonomous vehicle research, applying QAT to reduce computational costs.
2.3.3 Discussion
Tables 10 and 11 provide an overview of PTQ techniques for image classification and object detection on ViT models. A notable trend is the dominance of PTQ methods, with fewer studies exploring training-phase QAT. Table 10 highlights the top-1 accuracy improvements achieved by PTQ methods on ViT architectures for the classification tasks, while Table 11 summarizes the mean average precision (mAP) and APbox improvement achieved from the baseline techniques for the object detection task. Moreover, Table 12 summarizes QAT techniques for image classification and object detection. Interestingly, most QAT methods for ViTs, such as those leveraging KD [73, 84, 104, 105], focus on optimizing quantized models but lack experimental validation on edge devices. In contrast, approaches like PackQViT [109] have experimented with their QAT frameworks on mobile devices, pushing the boundaries of practical deployment. However, these quantization techniques have been broadly experimented on datasets like ImageNet-1k and COCO 2017, raising questions about their generalizability to specialized domains such as medical imaging or autonomous driving. This gap underscores the need for future research to explore versatile quantization strategies that cater to diverse application areas and resource-constrained edge devices.
It is important to note that the impact of model compression techniques such as pruning, quantization, and KD extends beyond theoretical efficiency. These methods directly influence how ViTs can be deployed on edge devices, especially when paired with hardware-aware acceleration strategies discussed in Section 4. Moreover, the practical effectiveness of these compression techniques often depends on the capabilities of software toolchains, which provide essential backend support for optimized deployment, as explored in Section 3.
Table 12: Results of different QAT techniques proposed for ViTs. The results of QD-BEV [72] are the mean average precision (mAP) rather than APbox.
Algorithm Code Key point Dataset Results
@@ 2.3.3 Discussion
Q-ViT  [73] Enables 3-bit ViT QAT. ImageNet-1k  [45] LSQ  [111] DeiT-B 4 4 83.0 ($+2.1$) -
@@
Q-DETR  [84] KD to enhance quantized DETR’s representation PASCAL VOC  [112] LSQ  [111] TR50 2 2 - 50.7 ($+8.1$)
SM-TR50 2 2 - 50.2 ($+7.9$)
COCO 2017  [106] TR50 4 4 - 39.4 ($+6.1$)
SM-TR50 4 4 - 38.3 ($+4.4$)
AQ-DETR  [104] QAT based on auxiliary queries for DETR PASCAL VOC  [112] Q-DETR  [84] TR50 4 4 - 53.7 ($+3.3$)
DE TR50 4 4 - 63.1 ($+2.0$)
COCO 2017  [106] TR50 4 4 - 40.2 ($+2.8$)
DE TR50 4 4 - 44.1 ($+3.4$)
PSAQ-ViT V2 [105] Data-free QAT using adaptive KD ImageNet-1k  [45] PSAQ-ViT  [113] DeiT-B 8 8 81.5 ($+2.4$) -
@@
COCO 2017  [106] Standard V2 DC 8 8 - 44.8 (+.$+.3$)
SC 8 8 - 50.9 ($+0.6$)
PackQViT  [109] 8-bit QAT for mobile devices ImageNet-1k  [45] Q-ViT  [73] DeiT-B 8 8 82.9 ($+0.5$) -
@@
COCO 2017
 [106]
PTQ
 [70]
TR50 8 8 - 60.0 ($-3.1$)
QD-BEV  [72] View-guided loss image and BEV features. NuScenes  [114]
BEV-DFQ
 [115]
BEV 8 8 - 40.6 ($+2.2$)
BEV-HAWQv3
 [116]
8 8 - 40.6 ($+3.0$)
BEV-PACT
 [117]
8 8 - 40.6 ($+3.2$)
  • 1.
    Note: Model names are shortened to save space in the table (SM TR50: SMCA-DeiT-R50).
    TR50: DETR-R50; DE: Deformable; DC: DeiT-S with Cascade Mask R-CNN;
    SC: Swin-S with Cascade Mask R-CNN; SM: SMCA;  BEV: BEVFormer.
@@ 3.1 Software Tools
Deploying deep learning models on heterogeneous platforms demands specialized software tools that bridge the gap between cutting-edge artificial intelligence (AI) research and real-world applications. These tools empower developers to optimize, accelerate, and seamlessly integrate AI models across different hardware architectures. Table 13 compares widely used deployment tools across four functional dimensions: model optimization (MO), model compilation (MC), high-level synthesis (HLS), and inference runtime (IR), with their hardware compatibility spanning FPGA, GPU, and CPU platforms.
As FPGAs offer highly parallel and reconfigurable hardware capabilities, deploying AI models on FPGAs requires specialized software tools for efficient hardware mapping, optimization, and deployment. Both Vivado Design Suite  [^3] Vivado design suite. Retrieved January 18, 2025, from https://www.amd.com/en/products/software/adaptive-socs-and-fpgas/vivado.html (Xilinx) and the Quartus Prime Design Software  [^4] Intel Quartus Prime. Retrieved January 18, 2025, from https://www.intel.com/content/www/us/en/products/details/fpga/development-tools/quartus-prime.html (Intel) offer advanced synthesis, converting high-level languages to hardware description language (HDL) and preoptimized AI accelerators IP cores (such as Xilinx DPU or Intel AI Suite) that help to accelerate the inference task.
Vitis AI  [^5] Vitis AI. Retrieved January 18, 2025, from https://www.xilinx.com/products/design-tools/vitis/vitis-ai.html (for Xilinx) and OpenVINO  [^6] Openvinotoolkit. Retrieved January 18, 2025, from https://github.com/openvinotoolkit/openvino (for Intel) provide support for partial model compression, hardware-optimized operator libraries, and integrated runtime environments. However, both frameworks are tightly coupled to their respective hardware ecosystems, which limits cross-platform portability. Each serves as a comprehensive deployment platform for FPGA-based applications, leveraging their corresponding hardware toolchains for synthesis, optimization, and implementation. Besides those two software tools, Hls4ml  [^7] Hls4ml. Retrieved January 18, 2025, from https://fastmachinelearning.org/hls4ml/index.html and FINN  [^8] FINN. Retrieved January 18, 2025, from https://xilinx.github.io/finn/ designed to explore deep neural network inference on FPGAs efficiently and swiftly. However, both these libraries are still in the experimental phase.
In the GPU domain, TensorRT  [^9] TensorRT. Retrieved January 18, 2025, from https://developer.nvidia.com/tensorrt uses as a popular inference optimizer and accelerating library for NVIDIA GPU-based edge devices. In a broader manner, Triton Inference Server  [^10] NVIDIA Triton Inference Server. Retrieved January 18, 2025, from https://github.com/triton-inference-server/server enables the scalable, production-ready serving of multiple deep learning models on NVIDIA GPUs, x86, and Arm CPUs, integrating seamlessly with frameworks like TensorFlow, Keras, and PyTorch. However, neither supports FPGAs. On the other hand, oneDNN 111111oneDNN. Retrieved January 18, 2025, from https://github.com/oneapi-src/oneDNN is a widely adopted, open-source performance library for deep learning applications, offering cross-platform support. It is highly optimized for Intel CPUs, integrated GPUs, and ARM-based processors, with ongoing experimental support for NVIDIA GPUs, AMD GPUs, and RISC-V architectures. ONNX Runtime  [^12] ONNX Runtime. Retrieved January 18, 2025, from https://onnxruntime.ai/ stands out as a versatile, hardware-agnostic inference engine, supporting deployment across CPUs, GPUs, and FPGAs via backend execution providers. It balances broad compatibility with modular optimization and is often preferred in cross-platform production environments.
In practice, the choice of tool depends on the target device, optimization needs, and ecosystem. For example, TensorRT and Triton Inference Server are ideal for GPU deployments, Vitis AI excels in customized FPGA acceleration, while ONNX Runtime is well-suited for flexible and portable edge inference across multiple hardware types. In conclusion, unified toolkits for multiple hardware platforms might be an interesting research direction in the future.
@@
Tool Type Key Features Supported Edge Devices
MO MC HLS IR FPGA GPU CPU
Xilinx Vivado design suite 3 Toolkit
@@
Intel Quartus Prime 4 Toolkit
@@
OpenVINO 6 Toolkit
@@
Vitis AI 5 Engine
@@
FINN 8 Engine
@@
TensorRT 9 Engine No
@@
ONNX Runtime 12 Engine Yes Yes Yes
Hls4ml 7 Library Yes No No
oneDNN 11 Library No Experimental Intel CPU, ARM
NVIDIA Triton
Inference Server 10
Inference Server No
@@
  • 1.
    MO: Model Optimization; MC: Model Compilation;
    HLS: High-Level Synthesis; IR: Inference Runtime